OpenStudy (anonymous):

Prove csc^2(theta) + cot^2 (theta) = 2csc^2(theta)-1

OpenStudy (anonymous):

cot^2(theta) = csc^2(theta)-1 csc^2(theta) + csc^2(theta) - 1 = 2csc^2(theta) - 1

OpenStudy (anonymous):

how did you go from cot^2 to csc^2 + csc^2 -1 in the next step

OpenStudy (anonymous):

The trig identity is cot^2(theta)+1 = csc^2(theta). I just switched that around for the first step.

OpenStudy (anonymous):

but it's just cot not cot +1

OpenStudy (anonymous):

csc^2 + cot^2 = 2csc^2 - 1 csc^2 + (csc^2-1) = 2csc^2-1 csc^2 + csc^2 -1 = 2csc^2-1

OpenStudy (anonymous):

ooooo ok thank you :D

OpenStudy (anonymous):

sorry about that